### v1

@@ 4.1 AI Generation Image Detection Benchmark
In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models. Therefore, we select only two commonly used public datasets: DiffusionForensics [59] and GenImage [69]. The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models. Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model [69]. The detailed data splits are summarized in Table D.1.3.
@@
ForensicHub implements five SoTA methods in AIGC detection: Dire [59], DualNet [63], HiFiNet [18], Synthbuster [2], and UnivFD [40], among which Synthbuster has no official open-source code and is fully reimplemented by us. More details about models and training settings can be found in Appendix E.1.
@@ Results.
Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics [59], and cross-domain results on different generative models and the total set on GenImage [69]. The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data. However, the relatively poor generalization to generative models like Midjourney and Wukong highlights areas for improvement and provides guidance for future model development.
@@ 4.2 Document Image Manipulation Localization Benchmark
DualNet [63](APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
HiFiNet [18](CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
Synthbuster(OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
UnivFD(CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
MVSS-Net(ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
CAT-Net(IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
Trufor(CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
IML-ViT(Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
Mesorch(AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
@@
Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE [58], OSTF [44], TPIC-13 [57], and RTM [32]; and sliced datasets, represented by Doctamper [43]. The primary difference lies in whether the images are preprocessed using patch-wise slicing.
@@
We employ two open-source models, DTD [43] and FFDN [6], and reproduce two closed-source models, CATFB [52] and TIFDM [12]. All details are provided in Appendix D.2.4 and E.2.
@@ Results.
Following the original protocols [43, 6], each detector is trained on its designated training split and evaluated on the corresponding test split. As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model. Notably, all three methods incorporate JPEG-specific priors, such as DCT coefficients and quantization tables, highlighting the discriminative value of compression artifacts for manipulation localization in document images.
Table 4: Binary-F1 scores of document detectors. AverageD is the mean over three Doctamper test dataset, and AverageAll is averaged across all seven test datasets.
@@
CAFTB [52](TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
DTD [43](CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
FFDN [6](ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
TIFDM [12](TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
MVSS-Net [5](ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
PSCC-Net [29](TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
Cat-Net [22](IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
IML-ViT [35](Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
Trufor [17](CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
Mesorch [70](AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
@@
CAFTB [52](TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
DTD [43](CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
FFDN [6](ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
TIFDM [12](TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
MVSS-Net [5](ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
PSCC-Net [29](TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
Cat-Net [22](IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
IML-ViT [35](Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
Trufor [17](CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
Mesorch [70](AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
@@ 5 Image Forensic Fusion Protocol
To explore the performance of different models under a unified forensic protocol, we implement an image forensic fusion protocol (IFF-Protocol) inspired by CAT-Net’s training data construction strategy. The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random. During training, we select FaceForensics++ [47] from Deepfake, CASIAv2 [11] from IMDL, GenImage [69] from AIGC, and OSTF [44], RealTextManipulation, T-SROIE, and Tampered-IC13 from the Document. We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch. During testing, we evaluate the models directly on datasets from different domains without fine-tuning.
@@
We resize images to 256×256 (except UnivFD, DTD and FFDN, see Appendix F for details) and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur. All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from $41e-$ 4 to $51e-$ 5. For models that output masks (IMDL and Doc), we apply max pooling to the final-layer feature maps to obtain a predicted label and compute the loss using only the label.
@@
Resnet [19] $44.55$ $20.59$ Capsule-Net [37] $3.90$ $32.43$ Mesorch [70] $85.75$ $57.95$
Xception [7] $22.86$ $12.05$ RECCE [4] $25.83$ $16.18$ DualNet [63] $7.99$ $66.34$
EfficientNet [55] $19.34$ $4.14$ SPSL [28] $20.81$ $12.06$ HiFiNet [18] $6.89$ $145.00$
ConvNeXt [31] $50.22$ $22.74$ MVSS-Net [5] $147.00$ $83.08$ UnivFD [40] $428.00$ $156.00$
Segformer [64] $44.07$ $16.29$ Trufor [17] $68.70$ $112.00$ DTD [43] $67.07$ $272.00$
Swin [30] $49.61$ $28.65$ IML-ViT [35] $91.78$ $80.37$ FFDN [6] $89.20$ $453.00$
@@
FF-c40 CDFv2 DFD Columbia IMD2020 Autosplice DF GenImage T-SROIE OSTF RTM
Resnet [19](CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
Xception [7](CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
EfficientNet [55](ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
Segformer [64](NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
Swin [30](ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
ConvNext [31](CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
Capsule-Net [37](ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
RECCE [4](CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
SPSL [28](CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
MVSS-Net [5](ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
Trufor [17](CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
IML-ViT [35](Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
Mesorch [70](AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
DualNet [63](APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
HiFiNet [18](CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
UnivFD [40](CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
DTD [43](CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
FFDN [6](ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
@@
Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection [15], DF refers to DiffusionForensics [59], and RTM refers to RealTextManipulation [32]. We provide detailed results in Appendix F.
The results show that surprisingly, visual backbones such as ConvNeXt [31] and Swin Transformer [30] outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images. Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks. For instance, UnivFD [40], a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 [39] from IMDL, revealing valuable insights into the transferability of cross-task methods.

### v2

@@ 4.1 AI Generation Image Detection Benchmark
In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models. Therefore, we select only two commonly used public datasets: DiffusionForensics wang2023dire and GenImage zhu2023genimage . The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models. Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model zhu2023genimage . The detailed data splits are summarized in Table D.1.3.
@@
ForensicHub implements five SoTA methods in AIGC detection: Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd , among which Synthbuster has no official open-source code and is fully reimplemented by us. More details about models and training settings can be found in Appendix E.1.
@@ Results.
Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics wang2023dire , and cross-domain results on different generative models and the total set on GenImage zhu2023genimage . The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data. However, the relatively poor generalization to generative models like Midjourney and Wukong highlights areas for improvement and provides guidance for future model development.
@@ 4.2 Document Image Manipulation Localization Benchmark
DualNet dualnet (APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
HiFiNet HiFi-Net2023 (CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
Synthbuster bammey2023synthbuster (OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
UnivFD univfd (CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
MVSS-Net MVSS_2021 (ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
CAT-Net CAT-Net2022 (IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
Trufor trufor2023 (CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
IML-ViT ma2023iml (Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
Mesorch zhu2025mesoscopic (AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
@@
Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , and RTM luo2025rtm ; and sliced datasets, represented by Doctamper qu2023doctamper . The primary difference lies in whether the images are preprocessed using patch-wise slicing.
@@
We employ two open-source models, DTD qu2023doctamper and FFDN chen2024ffdn , and reproduce two closed-source models, CATFB song2025caftb and TIFDM dong2024tifdm . All details are provided in Appendix D.2.4 and E.2.
@@ Results.
Following the original protocols qu2023doctamper ; chen2024ffdn , each detector is trained on its designated training split and evaluated on the corresponding test split. As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model. Notably, all three methods incorporate JPEG-specific priors, such as DCT coefficients and quantization tables, highlighting the discriminative value of compression artifacts for manipulation localization in document images.
Table 4: Binary‑F1 scores of document detectors. AverageD is the mean over three Doctamper test dataset, and AverageAll is averaged across all seven test datasets.
@@
CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
MVSS-Net MVSS_2021 (ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
@@
CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
MVSS-Net MVSS_2021 (ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
@@ 5 Image Forensic Fusion Protocol
To explore the performance of different models under a unified forensic protocol, we implement an image forensic fusion protocol (IFF-Protocol) inspired by CAT-Net’s training data construction strategy. The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random. During training, we select FaceForensics++ rossler2019faceforensics++ from Deepfake, CASIAv2 CASIA_2013 from IMDL, GenImage zhu2023genimage from AIGC, and OSTF qu2025ostf , RealTextManipulation luo2025rtm , T-SROIE wang2022tsroie , and Tampered-IC13 wang2022tpic from the Document. We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch. During testing, we evaluate the models directly on datasets from different domains without fine-tuning.
@@
We resize images to 256×256 (except UnivFD, DTD and FFDN, see Appendix F for details) and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur. All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from 1e41e-4 to 1e51e-5. For models that output masks (IMDL and Doc), we apply max pooling to the final-layer feature maps to obtain a predicted label and compute the loss using only the label.
@@
Resnet Resnet_2016 $44.55$ $20.59$ Capsule-Net nguyen2019capsule $3.90$ $32.43$ Mesorch zhu2025mesoscopic $85.75$ $57.95$
Xception chollet2017xception $22.86$ $12.05$ RECCE cao2022recce $25.83$ $16.18$ DualNet dualnet $7.99$ $66.34$
EfficientNet tan2019efficientnet $19.34$ $4.14$ SPSL liu2021spsl $20.81$ $12.06$ HiFiNet HiFi-Net2023 $6.89$ $145.00$
ConvNeXt Convnet_2022 $50.22$ $22.74$ MVSS-Net MVSS_2021 $147.00$ $83.08$ UnivFD univfd $428.00$ $156.00$
Segformer SegFormer_2021 $44.07$ $16.29$ Trufor trufor2023 $68.70$ $112.00$ DTD qu2023doctamper $67.07$ $272.00$
Swin Swin_2021 $49.61$ $28.65$ IML-ViT ma2023iml $91.78$ $80.37$ FFDN chen2024ffdn $89.20$ $453.00$
@@
$F$-c40 CDFv2 DFD Columbia IMD2020 Autosplice DF GenImage T-SROIE OSTF RTM
Resnet Resnet_2016 (CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
Xception chollet2017xception (CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
EfficientNet tan2019efficientnet (ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
Segformer SegFormer_2021 (NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
Swin Swin_2021 (ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
ConvNext Convnet_2022 (CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
Capsule-Net nguyen2019capsule (ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
RECCE cao2022recce (CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
SPSL liu2021spsl (CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
Sia sia (ECCV22) 0.629 0.584 0.671 0.653 0.483 0.626 0.593 0.748 0.610 0.677 0.574 0.622
Effort effort (ICML25) 0.805 0.846 0.930 0.979 0.861 0.943 0.930 0.992 0.960 0.834 0.732 0.892
MVSS-Net MVSS_2021 (ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
Trufor trufor2023 (CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
IML-ViT ma2023iml (Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
Mesorch zhu2025mesoscopic (AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
DualNet dualnet (APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
HiFiNet HiFi-Net2023 (CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
UnivFD univfd (CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
FatFormer fatformer (CVPR24) 0.842 0.770 0.866 0.199 0.585 0.784 0.941 0.999 0.983 0.806 0.751 0.758
CO-SPY co-spy (CVPR25) 0.819 0.780 0.875 0.460 0.716 0.779 0.940 0.989 0.969 0.836 0.748 0.829
DTD qu2023doctamper (CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
FFDN chen2024ffdn (ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
@@
Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection dfd2019 , DF refers to DiffusionForensics wang2023dire , and RTM refers to RealTextManipulation luo2025rtm . We provide detailed results in Appendix F.
The results show that surprisingly, visual backbones such as ConvNeXt Convnet_2022 and Swin Transformer Swin_2021 outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images. Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks. For instance, UnivFD univfd , a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 IMD20_2020 from IMDL, revealing valuable insights into the transferability of cross-task methods.